OpenStudy (anonymous):

What is the solution to the equation 9 –3x ≈ 7 ? Answer? x = 0.376 x = 0.295 x = –0.295 x = –0.376

OpenStudy (hoblos):

3x = 2 => x=2/3=0.67 but it is not any one of the options !!

OpenStudy (anonymous):

I'll make an educated guess with the info you gave, thanks!

OpenStudy (turingtest):

If it's\[9^{-3x}\approx 7\]then\[x\approx -0.295\]so I think that is the answer.

OpenStudy (turingtest):

@Swisha you should have written the problem as 9^(-3x)=7 so that it would be understood. Copy-paste is a bad idea for this reason.

OpenStudy (anonymous):

C